# Is DPO Superior to PPO for LLM Alignment? A Comprehensive Study

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is currently the most widely used method to align large language models (LLMs) with human preferences. Existing RLHF methods can be roughly categorized as either reward-based or reward-free. Novel applications such as ChatGPT and Claude leverage reward-based methods that first learn a reward model and apply actor-critic algorithms, such as Proximal Policy Optimization (PPO). However, in academic benchmarks, the state-of-the-art results are often achieved via reward-free methods, such as Direct Preference Optimization (DPO). Is DPO truly superior to PPO? Why does PPO perform poorly on these benchmarks? In this paper, we first conduct both theoretical and empirical studies on the algorithmic properties of DPO and show that DPO may have fundamental limitations. Moreover, we also comprehensively examine PPO and reveal the key factors for the best performances of PPO in fine-tuning LLMs. Finally, we benchmark DPO and PPO across various a collection of RLHF testbeds, ranging from dialogue to code generation. Experiment results demonstrate that PPO is able to surpass other alignment methods in all cases and achieve state-of-the-art results in challenging code competitions.
@@ 2 Related Work
This paper concentrates on RLHF methods, which can be broadly categorized into reward-based and reward-free approaches. Reward-based methods entail training a reward model on preference data in an initial phase (Gao et al., 2023; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). Subsequently, this learned reward model is utilized to provide a reward signal for online Reinforcement Learning (RL) algorithms such as PPO (Schulman et al., 2017). There exist previous works that have studied these methods through hyper-parameter tuning and analyzed the effects of the quality reward model quality (Zheng et al., 2023; Casper et al., 2023). In contrast, reward-free methods offer a simpler training procedure by directly training LLMs on preference data or ranking data to distill human preference (Yuan et al., 2023; Liu et al., 2023; Touvron et al., 2023; Rafailov et al., 2023; Song et al., 2023; Dong et al., 2023). Among these reward-free methods, DPO (Rafailov et al., 2023) has demonstrated strong performances and become popular in the community (MistralAI, 2023; Chen et al., 2024; Yuan et al., 2024). This paper examines representatives from both categories, specifically PPO and DPO. We analyze the limitations of DPO theoretically and empirically and explore the key factors for PPO training.
Concurrent efforts have been undertaken to avoid reward model overoptimization (Ramé et al., 2024), facilitate alignment data generation (Lee et al., 2023; Yang et al., 2023), and implement resource-efficient RLHF systems (Yao et al., 2023; Santacroce et al., 2023). These works complement our study and can be seamlessly integrated into our implementation. Previous works have explored the implementation details of PPO for LLMs (Zheng et al., 2023; Ramamurthy et al., 2023). Our paper extends its investigations with additional RLHF techniques, optimizing PPO performance to surpass its reward-free counterpart, DPO. This is an aspect not previously reported in the literature. Our work is also closely related to studies on algorithm implementation in the RL community (Engstrom et al., 2020; Andrychowicz et al., 2021; Yu et al., 2022). However, our findings provide further insights into fine-tuning LLMs with a model size of up to 34B parameters.
@@ 3 Preliminary
We remark that although Rafailov et al. (2023) performs a single-round DPO over the preference dataset, some recent works also adapt DPO to an iterative variant with a learned reward model (Yuan et al., 2024). We also investigate the performance of iterative DPO.
@@ 4.3 Experiments on Real Preference Datasets
We perform our experimental analysis on SafeRLHF dataset (Dai et al., 2023). In this dataset, preference pairs have the form $(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2},l_{\mathrm{h}},l_{\mathrm{s}},b_{1},b_{2})$, where $l_{\mathrm{h}}$ and $l_{\mathrm{s}}$ are a preference labels over $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ in terms of helpfulness and safety, respectively, which could be either 1 or 2. $b_{1}$ and $b_{2}$ are binary safety labels of these two responses. With this dataset, our objective is to train an LLM that prioritizes safety over helpfulness in content generation. Specifically, in constructing the preference dataset, our preference is for the more helpful response when both options are considered safe. However, if safety varies between the responses, our preference shifts towards the safer option. Following Dai et al. (2023), the base model is trained on the Alpaca (Taori et al., 2023) open-source dataset with SFT, denoted as SFT (Alpaca). We use the evaluation models released by the official codebase to evaluate the helpfulness and harmfulness. We remark that these official evaluation models are not involved during the PPO training. Our experimental study is shown in Table 2. Our observations are summarized below.

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is currently the most widely used method to align large language models (LLMs) with human preferences. Existing RLHF methods can be roughly categorized as either reward-based or reward-free. Novel applications such as ChatGPT and Claude leverage reward-based methods that first learn a reward model and apply actor-critic algorithms, such as Proximal Policy Optimization (PPO). However, in academic benchmarks, the state-of-the-art results are often achieved via reward-free methods, such as Direct Preference Optimization (DPO). Is DPO truly superior to PPO? Why does PPO perform poorly on these benchmarks? In this paper, we first conduct both theoretical and empirical studies on the algorithmic properties of DPO and show that DPO may have fundamental limitations. Moreover, we also comprehensively examine PPO and reveal the key factors for the best performances of PPO in fine-tuning LLMs. Finally, we benchmark DPO and PPO across a collection of RLHF testbeds, ranging from dialogue to code generation. Experiment results demonstrate that PPO is able to surpass other alignment methods in all cases and achieve state-of-the-art results in challenging code competitions.
@@ 2 Related Work
This paper concentrates on RLHF methods, which can be broadly categorized into reward-based and reward-free approaches. Reward-based methods entail training a reward model on preference data in an initial phase (Gao et al., 2023; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022). Subsequently, this learned reward model is utilized to provide a reward signal for online Reinforcement Learning (RL) algorithms such as PPO (Schulman et al., 2017). There exist previous works that have studied these methods through hyper-parameter tuning and analyzed the effects of the quality reward model quality (Zheng et al., 2023; Casper et al., 2023). In contrast, reward-free methods offer a simpler training procedure by directly training LLMs on preference data or ranking data to distill human preference (Yuan et al., 2023; Liu et al., 2023; Touvron et al., 2023; Rafailov et al., 2023; Song et al., 2023; Dong et al., 2023; Hong et al., 2024). Among these reward-free methods, DPO (Rafailov et al., 2023) has demonstrated strong performances and become popular in the community (MistralAI, 2023; Chen et al., 2024; Yuan et al., 2024). Recent work discussed the performance gap of DPO and PPO on synthetic contextual bandits (Li et al., 2023). In this paper, We analyze the limitations of DPO theoretically and empirically, and explore the key factors for PPO training.
Concurrent efforts have been undertaken to avoid reward model overoptimization (Ramé et al., 2024), facilitate alignment data generation (Lee et al., 2023; Yang et al., 2023), and implement resource-efficient RLHF systems (Yao et al., 2023; Santacroce et al., 2023). These works complement our study and can be seamlessly integrated into our implementation. Previous works have explored the implementation details of PPO for LLMs (Zheng et al., 2023; Ramamurthy et al., 2023). Our paper extends its investigations with additional RLHF techniques, optimizing PPO performance to surpass its reward-free counterpart, DPO. Our work is also closely related to studies on algorithm implementation in the RL community (Engstrom et al., 2020; Andrychowicz et al., 2021; Yu et al., 2022). However, our findings provide further insights into fine-tuning LLMs with a model size of up to 34B parameters.
@@ 3 Preliminary
We remark that although Rafailov et al. (2023) performs a single-round DPO over the preference dataset, some recent works also adapt DPO to an iterative variant with a learned reward model (Xiong et al., 2023; Yuan et al., 2024). We also investigate the performance of iterative DPO.
@@ 4.3 Experiments on Real Preference Datasets
We perform our experimental analysis on SafeRLHF dataset (Dai et al., 2023). In this dataset, preference pairs have the form $(\mathbf{x},\mathbf{y}_{1},\mathbf{y}_{2},l_{\mathrm{h}},l_{\mathrm{s}},b_{1},b_{2})$, where $l_{\mathrm{h}}$ and $l_{\mathrm{s}}$ are a preference labels over $\mathbf{y}_{1}$ and $\mathbf{y}_{2}$ in terms of helpfulness and safety, respectively, which could be either 1 or 2. $b_{1}$ and $b_{2}$ are binary safety labels of these two responses, which could be positive or negative. With this dataset, our objective is to train an LLM that prioritizes safety over helpfulness in content generation. Specifically, in constructing the preference dataset, our preference is for the more helpful response when both options are considered safe (i.e., $l_{h}$ if $b_{1}$ and $b_{2}$ are both positive). Otherwise, our preference shifts towards the safer one (i.e., $l_{s}$). Following Dai et al. (2023), the base model is trained on the Alpaca (Taori et al., 2023) open-source dataset with SFT, denoted as SFT (Alpaca). We use the evaluation models released by the official codebase to evaluate the helpfulness and harmfulness. We remark that these official evaluation models are not involved during the training. Our experimental study is shown in Table 2.